# Automatically Interpreting Millions of Features in Large Language Models

### v2

@@ Abstract
While the activations of neurons in deep neural networks usually do not have a simple human-understandable interpretation, sparse autoencoders (SAEs) can be used to transform these activations into a higher-dimensional latent space which can be more easily interpretable. However, SAEs can have millions of distinct latents, making it infeasible for humans to manually interpret each one. In this work, we build an open-source automated pipeline to generate and evaluate natural language interpretations for SAE latents using LLMs. We test our framework on SAEs of varying sizes, activation functions, and losses, trained on two different open-weight LLMs. We introduce five new techniques to score the quality of interpretations that are cheaper to run than the previous state of the art. One of these techniques, intervention scoring, evaluates the interpretability of the effects of intervening on a latent, which we find explains latents that are not recalled by existing methods. We propose guidelines for generating better interpretations that remain valid for a broader set of activating contexts, and discuss pitfalls with existing scoring techniques. Our code is available here, and our interpretations are available here.
@@ 1 Introduction
Large language models (LLMs) have reached human level performance in a broad range of domains (OpenAI, 2023), and can even be leveraged to develop agents (Wang et al., 2023) that can strategize (Bakhtin et al., 2022), cooperate and develop new ideas (Lu et al., 2024; Shaham et al., 2024). At the same time, we understand little about the internal representations driving their behavior. Early mechanistic interpretability research focused on analyzing the activation patterns of individual neurons (Olah et al., 2020; Gurnee et al., 2023; 2024). Due to the large number of neurons to be interpreted, automated approaches were proposed (Bills et al., 2023), where a second LLM is used to propose an interpretation given a set of neuron activations and the text snippets it activates on, in a process similar to that of generating a human label. But research has found that most neurons are “polysemantic”, activating in contexts that can be significantly different (Arora et al., 2018; Elhage et al., 2022). The Linear Representation Hypothesis (Park et al., 2023) posits that human-interpretable concepts are encoded in linear combinations of neurons. A significant branch of current interpretability work focuses on extracting these latents and disentangling them (Bereska & Gavves, 2024).
Sparse autoencoders (SAEs) were proposed as a way to address polysemanticity (Cunningham et al., 2023). SAEs consist of two parts: an encoder that transforms activation vectors into a sparse, higher-dimensional latent space, and a decoder that projects the latents back into the original space. Both parts are trained jointly to minimize reconstruction error. SAE latents were found to be interpretable and potentially more monosemantic than neurons (Bricken et al., 2023; Cunningham et al., 2023). Recently, a significant effort was made to scale SAE training to larger models, like GPT-4 (Gao et al., 2024) and Claude 3 Sonnet (Templeton et al., 2024), and they have become an important interpretability tool for LLMs.
Training an SAE yields numerous sparse features, each of which needs a natural language interpretation. In this work, we introduce an automated framework that uses LLMs to generate an interpretation for each latent in an SAE. We use this framework to explain millions of latents across multiple models, layers, and SAE architectures. We also propose new ways to evaluate the quality of interpretations, and discuss the problems with existing approaches. We hope that these result can inform other repositories of interpretations for SAE latents that already exist (Lin & Bloom, 2023), as well as improve comparisons between different SAE architectures.
@@ 2 Related work
One of the most successful approaches to automated interpretability focused on explaining neurons of GPT-2 using GPT-4 (Bills et al., 2023). GPT-4 was shown examples of contexts where a given neuron was active and was tasked to provide a short interpretation that could capture the activation patterns. To evaluate if a given interpretation captured the behavior of the neuron, GPT-4 was tasked to predict the activations of the neuron in a given context having access to that interpretation. The interpretation is then scored by how much the simulated activations correlate with the true activations. A similar approach was used in Templeton et al. (2024) to explain SAE latents of Claude 3 Sonnet. In general, current approaches focus on collecting contexts together with latent activations from the model to be explained, and use a larger model to find patterns in activating contexts.
Following those works, other methods of evaluating interpretations have been proposed, including asking a model to generate activating examples and measuring how much a ”module” activates (Singh et al., 2023; Kopf et al., 2024). More recently, “interpretability agents” have been built, managing to do iterative experiment to find the best interpretations of vision neurons (Shaham et al., 2024). Potentially cheaper versions of automated interpretability have been suggested, where the model that is being explained doubles as an interpretation generation model (Kharlapenko et al., 2024). A prompt querying the meaning of a single placeholder token is passed to the model and latent activations are patched into its residual stream at the position of the placeholder token during execution, generating continuations related to the latent. This technique is inspired by earlier work on Patchscopes (Ghandeharioun et al., 2024) and SelfIE (Chen et al., 2024).
@@ 3 Automated interpretability pipeline
In this Section we explain, step-by-step, the main pipeline used to generate interpretations and score them. The pipeline can be broadly divided into 3 sequential steps. First, the activations of the SAEs to be interpreted are collected over a broad dataset. Then, for each latent, relevant contexts are selected and shown to an LLM which generates an interpretation for the activation pattern observed. Finally, these interpretations are matched with different contexts and used by an LLM in different tasks that evaluate how good the interpretations are in predicting activating and non-activating contexts, details explained below. As an illustration, we represent how that pipeline might look like for a real latent, see Figure 1.
Figure 1: Auto-interpretability pipeline. The first step of the interpretability pipeline is collecting the activations of the target SAE over a broad range of text. In this figure we show short contexts for legibility, and over represent the active latent, which in reality would be active in a very small portion of the text. The activating contexts are selected and shown to an explainer LLM, prompt in Appendix A.2, which provides a short interpretation. This interpretation is then given to a scorer LLM that is tasked to use this interpretation to distinguish activating from non activating contexts, see Section 3.3
@@ 3.1 Collecting activations
Sparse autoencoders (SAEs) are feedforward networks with one hidden layer trained to reproduce their input using a sparse number of neurons. This means that while there are several times more neurons than there are dimensions of the input, only a small fraction of these neurons– less than 60 in this work– are non-zero and contribute to the output. In our pipeline, we seek to explain the non-zero activations of the SAE hidden layer neurons, which we call “latents.”
We collected latent activations from the SAEs over a 10M token sample of RedPajama-v2 (RPJv2;  Computer 2023). This is the same dataset that we used to train the Llama 3.1 8b SAEs, and consists of a data mix similar to the Llama 1 pretraining corpus.
We collected batches of 256 tokens starting with the beginning of sentence token (BOS). [^1] We throw out the activations on the BOS token when generating interpretations for Gemma, as we were told in personal communication that these SAEs were not trained on BOS activations. The contexts used for activation collection are smaller than the contexts used to train the SAEs, and we find that on average, 30% of the latents of the 131k latent, per layer, Gemma 2 9b SAE don’t activate more than 200 times over these 10M tokens and 15% don’t activate at all. When we consider the training context length of 1024, only 5% of latents don’t fire. When using a closer proxy of the training data, the “un-copyrighted” Pile, we find that the number of latents that activate fewer than 200 times decreases to 15%, and that only 1% don’t activate at all, even when considering contexts of size 256. Interestingly, in the case of the 16k latent Gemma 2 9b SAE only around 10% of latents activate fewer than 200 times on the RPJv2 sample, suggesting that the larger 131k latent SAE learned more dataset specific latents than its smaller cousin.
@@ 3.2 Generating interpretations
Our approach to generating interpretations follows Bricken et al. (2023) in showing the explainer model examples sampled from different quantiles, but uses a more “natural” prompt where the activating example is shown whole, with the activating tokens emphasized and their activation strength shown after the example, see Appendix A.2 for the full prompt. For examples of some interpretations and their scores, see Figure A1.
We show the explainer model, Llama 3.1 70b Instruct, 40 different activating examples, each activating example consisting of 32 tokens, where the activating tokens can be on any position. While most latents are well-explained by contexts of length 32, we expect latents active over longer contexts to not be well captured by our pipeline. Using longer contexts to generate explanations would require either showing the model less examples, or using a much more powerful (and expensive) explainer model. Because we are able to evaluate interpretations significantly cheaper, bad interpretation can easily be filtered out, and we leave the explanation of “long-range” latents for future work. Using this interpretation setup would cost \$200 to interpret 1 million latents.
We find that randomly sampling from a broader set of examples leads to interpretations that cover a more diverse set of activating examples, sometimes to the detriment of the top activating examples, see Figure 3. Using only top activating examples often yields more concise and specific interpretations that accurately describe these examples, but which fail to capture the whole distribution. On the other hand, stratified sampling from the deciles of the activation distribution can lead to interpretations that are too broad and that fail to capture a meaningful interpretation. For examples of these types of failure modes, see Figure A2 and the discussion in the Appendix A.3.
@@ 3.2.1 Automatically interpreting interventions
Automatic interpretability methods, including most of those explored in this paper, typically look for correlations between activation of a latent and some natural-language property of the input. However, some latents are more closely related to what the model will output. For example, we find a latent222latent 157 of Gemma 2 9b’s layer 32 autoencoder with 131k latents and an average L0 norm of 51. It has a detection score of 0.6. whose activation causes the model to output words associated with reputation but does not have a simple interpretation in terms of inputs.
We define an output latent as a latent that causes a property of the model’s output that can be easily explained in natural language. See Section 3.3.5 for the definition we use in scoring.
Output latents can also be described in terms of their correlation with inputs. For example, the “reputation” latent activates in contexts where likely next tokens relate to reputation. However, explaining output latents in terms of causal influence on output has two advantages.
@@
    Scalability. Output latents are easier to describe in terms of effect on output because the explainer only needs to notice a simple pattern of output. The pattern of inputs this latent correlates with is more complex. Explaining a latent by correlating it with inputs requires approximating the computation of the subject model leading up to that latent, which may be challenging when subject models are highly capable and performing difficult tasks. Some latents’ influence on output, however, might remain easily explainable.
@@
    Causal evidence. We might like to know that a latent causes a property of the model’s output so that we can steer the model. Previous work has shown that existing auto-interpretability scores fail to accurately capture how well a given interpretation can predict the effect of intervening on a given neuron (Huang et al., 2023). Further, prior work has argued that causal evidence is more robust to distribution shifts (Bühlmann, 2018; Schölkopf et al., 2012).
@@ 3.3 Scoring interpretations
Not only should the generation of interpretations efficient, but so should the scoring of such interpretations. Practitioners would like to know how faithful each interpretation is to the actual behavior of the network. Some latents may not be amenable to a simple interpretation, and in these cases we expect the auto-interpretability pipeline to output an interpretation with poor evaluation metrics. Secondly, we can use evaluations as a feedback signal for the training of explainer models and tuning hyperparameters in the pipeline. Finally, some SAEs may simply be poor quality overall, and the aggregate evaluation metrics of the SAE’s interpretations can be used to detect this.
@@
The quality of SAE interpretations has so far been measured via simulation scoring, a technique introduced by Bills et al. (2023) for evaluating interpretations of neurons. It involves asking an explainer model to “simulate” a latent in a set of contexts, which means predicting how strongly the latent should activate at each token in a context given an interpretation. The Pearson correlation between the simulated and true activations is the simulation score. The standard in the literature is to sample contexts from a “top-and-random” distribution that mixes maximally activating contexts and contexts sampled uniformly at random from the activating corpus. While oversampling the top activating contexts introduces bias, it is used as a cheap variance reduction technique, given that simulation scoring over hundreds of examples per latent would be expensive (Cunningham et al., 2023).
@@
In this work, we take a slightly different view of what makes for a good interpretation: an interpretation should serve as a binary classifier distinguishing activating from non-activating contexts. The reasoning behind this is simple. Given the highly sparse nature of SAE latents, most of their variance could be captured by a binary predictor that predicts the mean nonzero activation when the latent is expected to be active, and zero otherwise. In statistics, zero-inflated phenomena are often modeled as a mixture distribution with two components: a Dirac delta on zero, and another distribution (e.g. Poisson) for nonzero values. This goes against the general wisdom of simulation scoring, which focuses only on activating examples, even though they are less than 0.01% of the relevant contexts for each latent of an SAE and that being able to distinguish non-activating contexts from activating contexts seems to be relatively more important.
Figure 2: The new proposed scoring methods. In detection scoring, the scorer model is tasked with selecting the set of sentences that activate a given latent given an interpretation. In this work, we show 5 examples at the same time, and each has an identical probability of being a sentence that activates the latent, independent of whether any other example also activates the latent. The activating tokens are colored in green for display, but that information is not shown to the scorer model. For fuzzing scoring, the scorer model is tasked with selecting the sentences where the highlighted tokens are the tokens that activate a target latent given an interpretation of that latent. In surprisal scoring, activating and non-activating examples are run through the model and the loss over those sentences is computed. Correct interpretations should decrease the loss in activating sentences compared to a generic interpretation, but shouldn’t significantly decrease the loss in non-activating sentences. For embedding scoring, activating and non-activating sentences are embedded as “documents” that should be retrieved using the interpretation as a query.
@@
As an alternative to simulation scoring, we introduce four new evaluation methods that focus on how well an interpretation enables a scorer to discriminate between activating and non-activating contexts. As an added benefit, all of these methods are more compute-efficient than simulation, see Table 1. We also propose a scoring method that evaluates the interpretability of interventions using a specific latent.

### v3

@@ Abstract
While the activations of neurons in deep neural networks usually do not have a simple human-understandable interpretation, sparse autoencoders (SAEs) can be used to transform these activations into a higher-dimensional latent space which can be more easily interpretable. However, SAEs can have millions of distinct features, making it infeasible for humans to manually interpret each one. In this work, we build an open-source automated pipeline to generate and evaluate natural language interpretations for SAE features using LLMs. We test our framework on SAEs of varying sizes, activation functions, and losses, trained on two different open-weight LLMs. We introduce five new techniques to score the quality of interpretations that are cheaper to run than the previous state of the art. One of these techniques, intervention scoring, evaluates the interpretability of the effects of intervening on a feature, which we find explains features that are not recalled by existing methods. We propose guidelines for generating better interpretations that remain valid for a broader set of activating contexts, and discuss pitfalls with existing scoring techniques. Our code is available online.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have reached human level performance in a broad range of domains (OpenAI, 2023), and can even be leveraged to develop agents (Wang et al., 2023) that can strategize (Bakhtin et al., 2022), cooperate and develop new ideas (Lu et al., 2024; Shaham et al., 2024). At the same time, we understand little about the internal representations driving their behavior. Early mechanistic interpretability research focused on analyzing the activation patterns of individual neurons (Olah et al., 2020; Gurnee et al., 2023, 2024). Due to the large number of neurons to be interpreted, automated approaches were proposed (Bills et al., 2023), where a second LLM is used to propose an interpretation given a set of neuron activations and the text snippets it activates on, in a process similar to that of generating a human label. But research has found that most neurons are “polysemantic”, activating in contexts that can be significantly different (Arora et al., 2018; Elhage et al., 2022). The Linear Representation Hypothesis (Park et al., 2023) posits that human-interpretable concepts are encoded in linear combinations of neurons. A significant branch of current interpretability work focuses on extracting these features and disentangling them (Bereska & Gavves, 2024).
Sparse autoencoders (SAEs) were proposed as a way to address polysemanticity (Cunningham et al., 2023). SAEs consist of two parts: an encoder that transforms activation vectors into a sparse, higher-dimensional latent space, and a decoder that projects the features back into the original space. Both parts are trained jointly to minimize reconstruction error. SAE features [^1] We use feature to mean the intermediate dimension of the SAE like in (Bricken et al., 2023). were found to be interpretable and potentially more monosemantic than neurons (Bricken et al., 2023; Cunningham et al., 2023). Recently, a significant effort was made to scale SAE training to larger models, like GPT-4 (Gao et al., 2024) and Claude 3 Sonnet (Templeton et al., 2024), and they have become an important interpretability tool for LLMs.
Training an SAE yields numerous sparse features, each of which needs a natural language interpretation. In this work, we introduce an automated framework that uses LLMs to generate an interpretation for each feature in an SAE. We use this framework to explain millions of features across multiple models, layers, and SAE architectures. We also propose new ways to evaluate the quality of interpretations, and discuss the problems with existing approaches. We hope that these result can inform other repositories of interpretations for SAE features that already exist (Lin & Bloom, 2023), as well as improve comparisons between different SAE architectures.
@@ 2 Related Work
One of the most successful approaches to automated interpretability focused on explaining neurons of GPT-2 using GPT-4 (Bills et al., 2023). GPT-4 was shown examples of contexts where a given neuron was active and was tasked to provide a short interpretation that could capture the activation patterns. To evaluate if a given interpretation captured the behavior of the neuron, GPT-4 was tasked to predict the activations of the neuron in a given context having access to that interpretation. The interpretation is then scored by how much the simulated activations correlate with the true activations. A similar approach was used in Templeton et al. (2024) to explain SAE features of Claude 3 Sonnet. In general, current approaches focus on collecting contexts together with feature activations from the model to be explained, and use a larger model to find patterns in activating contexts.
Following those works, other methods of evaluating interpretations have been proposed, including asking a model to generate activating examples and measuring how much a ”module” activates (Singh et al., 2023; Kopf et al., 2024). More recently, “interpretability agents” have been built, managing to do iterative experiment to find the best interpretations of vision neurons (Shaham et al., 2024). Potentially cheaper versions of automated interpretability have been suggested, where the model that is being explained doubles as an interpretation generation model (Kharlapenko et al., 2024). A prompt querying the meaning of a single placeholder token is passed to the model and feature activations are patched into its residual stream at the position of the placeholder token during execution, generating continuations related to the feature. This technique is inspired by earlier work on Patchscopes (Ghandeharioun et al., 2024) and SelfIE (Chen et al., 2024).
@@ 3 Automated interpretability pipeline
In this section we explain, step-by-step, the main pipeline used to generate interpretations and score them. The pipeline can be broadly divided into 3 sequential steps. First, the activations of the SAEs to be interpreted are collected over a broad dataset. Then, for each feature, relevant contexts are selected and shown to an LLM which generates an interpretation for the activation pattern observed. Finally, these interpretations are matched with different contexts and used by an LLM in different tasks that evaluate how good the interpretations are in predicting activating and non-activating contexts, details explained below. As an illustration, we represent how that pipeline might look like for a real feature, see Figure 1.
Figure 1: Auto-interpretability pipeline. The first step of the interpretability pipeline is collecting the activations of the target SAE over a broad range of text. In this figure we show short contexts for legibility, and over represent the active feature, which in reality would be active in a very small portion of the text. The activating contexts are selected and shown to an explainer LLM, prompt in Appendix B, which provides a short interpretation. This interpretation is then given to a scorer LLM that is tasked to use this interpretation to distinguish activating from non activating contexts, see Section 3.3
@@ 3.1 Collecting activations
Sparse autoencoders (SAEs) are feedforward networks with one hidden layer trained to reproduce their input using a sparse number of neurons. This means that while there are several times more neurons than there are dimensions of the input, only a small fraction of these neurons– less than 60 in this work– are non-zero and contribute to the output. In our pipeline, we seek to explain the non-zero activations of the SAE hidden layer neurons, which we call “features.”
We collected feature activations from the SAEs over a 10M token sample of RedPajama-v2 (RPJv2;  Computer 2023). This is the same dataset that we used to train the Llama 3.1 8b SAEs, and consists of a data mix similar to the Llama 1 pretraining corpus.
We collected batches of 256 tokens starting with the beginning of sentence token (BOS). [^2] We throw out the activations on the BOS token when generating interpretations for Gemma, as we were told in personal communication these SAEs were not trained on BOS activations. The contexts used for activation collection are smaller than the contexts used to train the SAEs, and we find that, on average, 30% of the features, per layer, of the 131k feature Gemma 2 9b SAE don’t activate more than 200 times over these 10M tokens and 15% don’t activate at all. When we consider the training context length of 1024, only 5% of features don’t fire. When using a closer proxy of the training data, the “un-copyrighted” Pile, we find that the number of features that activate fewer than 200 times decreases to 15%, and that only 1% don’t activate at all, even when considering contexts of size 256. Interestingly, in the case of the 16k feature Gemma 2 9b SAE only around 10% of features activate fewer than 200 times on the RPJv2 sample, suggesting that the larger 131k feature SAE learned more dataset specific features than its smaller cousin.
@@ 3.2 Generating interpretations
Our approach to generating interpretations follows Bricken et al. (2023) in showing the explainer model examples sampled from different quantiles, but uses a more “natural” prompt where the activating example is shown whole, with the activating tokens emphasized and their activation strength shown after the example, see Appendix B for the full prompt, similar to what is done in (Lee et al., 2023). For examples of some interpretations and their scores, see Figure A1.
We show the explainer model, Llama 3.1 70b Instruct, 40 different activating examples, each activating example consisting of 32 tokens, where the activating tokens can be on any position. While most features are well-explained by contexts of length 32, we expect features active over longer contexts to not be well captured by our pipeline. Using longer contexts to generate explanations would require either showing the model less examples, or using a much more powerful (and expensive) explainer model. Because we are able to evaluate interpretations significantly cheaper, bad interpretation can easily be filtered out, and we leave the explanation of “long-range” features for future work. Using this interpretation setup would cost \$200 to interpret 1 million features.
We find that randomly sampling from a broader set of examples leads to interpretations that cover a more diverse set of activating examples, sometimes to the detriment of the top activating examples, see Figure 2. Using only top activating examples often yields more concise and specific interpretations that accurately describe these examples, but which fail to capture the whole distribution. On the other hand, stratified sampling from the deciles of the activation distribution can lead to interpretations that are too broad and that fail to capture a meaningful interpretation. For examples of these types of failure modes, see Figure A2 and the discussion in the Appendix C.
@@ 3.2.1 Automatically interpreting interventions
Automatic interpretability methods, including most of those explored in this paper, typically look for correlations between activation of a feature and some natural-language property of the input. However, some features are more closely related to what the model will output. For example, we find a feature333feature 157 of Gemma 2 9b’s layer 32 autoencoder with 131k features and an average L0 norm of 51. Its detection score is 0.6. whose activation causes the model to output words associated with reputation but does not have a simple interpretation in terms of inputs.
We define an output feature as a feature that causes a property of the model’s output that can be easily explained in natural language. See Section 3.3.5 for the definition we use in scoring.
Output features can also be described in terms of their correlation with inputs. For example, the “reputation” feature activates in contexts where likely next tokens relate to reputation. However, explaining output features in terms of causal influence on output has two advantages.
@@
    Scalability. Output features are easier to describe in terms of effect on output because the explainer only needs to notice a simple pattern of output. The pattern of inputs this feature correlates with is more complex. Explaining a feature by correlating it with inputs requires approximating the computation of the subject model leading up to that feature, which may be challenging when subject models are highly capable and performing difficult tasks. Some features’ influence on output, however, might remain easily explainable.
@@
    Causal evidence. We might like to know that a feature causes a property of the model’s output so that we can steer the model. Previous work has shown that existing auto-interpretability scores fail to accurately capture how well a given interpretation can predict the effect of intervening on a given neuron (Huang et al., 2023). Further, prior work has argued that causal evidence is more robust to distribution shifts (Bühlmann, 2018; Schölkopf et al., 2012).
@@ 3.3 Scoring interpretations
Not only should the generation of interpretations efficient, but so should the scoring of such interpretations. Practitioners would like to know how faithful each interpretation is to the actual behavior of the network. Some features may not be amenable to a simple interpretation, and in these cases we expect the auto-interpretability pipeline to output an interpretation with poor evaluation metrics. Secondly, we can use evaluations as a feedback signal for the training of explainer models and tuning hyperparameters in the pipeline. Finally, some SAEs may simply be poor quality overall, and the aggregate evaluation metrics of the SAE’s interpretations can be used to detect this.
@@
The quality of SAE interpretations has so far been measured via simulation scoring, a technique introduced by Bills et al. (2023) for evaluating interpretations of neurons. It involves asking an explainer model to “simulate” a feature in a set of contexts, which means predicting how strongly the feature should activate at each token in a context given an interpretation. The Pearson correlation between the simulated and true activations is the simulation score. The standard in the literature is to sample contexts from a “top-and-random” distribution that mixes maximally activating contexts and contexts sampled uniformly at random from the activating corpus. While oversampling the top activating contexts introduces bias, it is used as a cheap variance reduction technique, given that simulation scoring over hundreds of examples per feature would be expensive (Cunningham et al., 2023).
@@
In this work, we take a slightly different view of what makes for a good interpretation: an interpretation should serve as a binary classifier distinguishing activating from non-activating contexts. The reasoning behind this is simple. Given the highly sparse nature of SAE features, most of their variance could be captured by a binary predictor that predicts the mean nonzero activation when the feature is expected to be active, and zero otherwise. In statistics, zero-inflated phenomena are often modeled as a mixture distribution with two components: a Dirac delta on zero, and another distribution (e.g. Poisson) for nonzero values. This goes against the general wisdom of simulation scoring, which focuses only on activating examples, even though they are less than 0.01% of the relevant contexts for each feature of an SAE and that being able to distinguish non-activating contexts from activating contexts seems to be relatively more important.
@@
As an alternative to simulation scoring, we introduce four new evaluation methods that focus on how well an interpretation enables a scorer to discriminate between activating and non-activating contexts. As an added benefit, all of these methods are more compute-efficient than simulation, see Table 1. We also propose a scoring method that evaluates the interpretability of interventions using a specific feature.